Model A takes 2hours to assemble and cost $25 per bike, to make.
Model B takes 3hours to assemble and cost $30 per bike, to make.
A company has a total of 34 hours and $350 per day to make models.
How many of each model can be made in a day? Solve equation using the substitution method or the elimination method

Answer by mananth(16949)

Model A = x numbers
Model B = y numbers
cost equation
25x+30y=350.............1
time equation
2x+3y=34............2
multiply equation 2 by -10 and add to equation 1
-20x-30y+25x+30y=-340+350
5x=10
x= 2 Model A
..
2*2+3y=34
4+3y=34
3y=34-4
3y=30
y=10 Model B